A ball thrown vertically upward reaches a maximum height of 50m above the level of projection. Calculate the

  1. time taken to reach maximum height
  2. Speed of the throw
1
Expert's answer
2021-03-10T17:15:58-0500

The speed of the throw can be found from the energy conservation law: the initial kinetic energy of the ball is equal to its potential energy.


"\dfrac{mv^2}{2} = mgh\\"

where "m" is the mass of the ball, "v" is its initial speed, "h = 50m" is the maximum height, "g = 9.8N\/kg" is the gravitational acceleration. Expressing the speed, obtain:


"v = \sqrt{2gh}\\
v = \sqrt{2\cdot 9.8\cdot 50} \approx 31.3m/s"

The distance covered by the ball is given by the kinematic equation:


"h = vt - \dfrac{gt^2}{2}\\
\dfrac{gt^2}{2} - \sqrt{2gh}t+h=0"

where "t" is the time.

Solving the quadratic equation, obtain:


"D = 2gh-2gh = 0\\
t = \dfrac{\sqrt{2gh}}{g} = \sqrt{\dfrac{2h}{g}}"


When "h = 50m", the time will be:


"t = \sqrt{\dfrac{2\cdot 50}{9.8}} \approx 5.1s"

Answer. 1) 5.1s, 2) 31.3 m/s.